The vapour pressure of a pure liquid air is 40 mm of HG at 310 Kelvin . the vapour pressure of this liquid in a solution with liquid b is 32 mm HG. mole fraction of a in the solution if it obays Raults lawis

Grade:12

1 Answers

ankur sharma
6 years ago
we know that, p​s=p*x​a where p = pressure on mixing solute and p*=pressure of pure solvent
and x​a= mole fraction of solvent N= moles of solvent n =moles of solute xb​=mole fraction of solute
      p​s = p*x​a   , dividing both side with p*
     p​s / p* = x​​a   substracting  from 1 both side 1- p/p*=  1-xa
        p*-p​s/p* = x​b  ​(x​a + x​b =1)   so by given conditions , p*=40 p​s​=32 and x​b=?
      40-32/40 = xb  so xb= 1/5